OpenStudy (cyclone3260):

Which of the following equations is of a parabola with a vertex at (0, -5)?

OpenStudy (cyclone3260):

a)y = (x - 5)2 b)y = (x + 5)2 c)y = x2 - 5 or d)y = x2 + 5

OpenStudy (anonymous):

Do you know this conics formula? y=a (x-h)^2 + k

OpenStudy (cyclone3260):

oh yea so d? @study100

OpenStudy (xapproachesinfinity):

the vertex is (h,k) in our case (0,-5) plug this into the formula y=a(x-0)^2-5 then y=ax^2-5 in your question they didn't consider a so it should be y=x^2-5